OpenStudy (anonymous):

Find the area of the region bounded by the curve y=x^3/1000 and the x-axis from x=0 to x=1 Please help?

geerky42 (geerky42):

That would be \(\displaystyle \int_0^1 \dfrac{x^3}{1000} dx\)

OpenStudy (anonymous):

just to update... if you need them, the answer choices are 0.00015 0.0002 0.00025 0.00033

OpenStudy (anonymous):

@geerky42 When I distribute the x over to the equation, do I add 3+1 to 1000?

geerky42 (geerky42):

multiply

geerky42 (geerky42):

(3+1)1000

geerky42 (geerky42):

to make anything easier, you can take 1/1000 out of integral, like this: \[\large \int_0^1 \dfrac{x^3}{1000}dx = \dfrac{1}{1000}\int_0^1x^3dx\]

geerky42 (geerky42):

So can you figure out the answer?

OpenStudy (shamim):